# PowerInfer: Fast Large Language Model Serving with a Consumer-grade GPU

### v1

@@ 8.3.1 Performance Breakdown
Figure 15 breaks down the contributions of each PowerInfer component to the overall performance speedup. Using a step-by-step integration method, we progressively incorporate PowerInfer features into llama.cpp. First, we add PowerInfer’s predictors and neuron-aware operators into llama.cpp (labeled "+PO"), enabling computation of only activated neurons on both GPU and CPU. Yet, +PO still adheres to layer-wise computation, where each layer is processed entirely by either GPU or CPU.
[…]
@@
The initial integration of +PO into llama.cpp yields performance boosts of 1.98$\times$ and 2.00$\times$ for OPT-30B and OPT-66B, respectively, primarily by reducing unnecessary inactive neurons. +Engine further escalates these gains to 9.97$\times$ and 3.43$\times$, thanks to precise neuron placement and intra-layer calculations that significantly increase the GPU’s computational share. Finally, incorporating +Policy results in improvements of 10.47$\times$ and 3.67$\times$. The enhancement achieved by our policy lies in its ability to finely balance the intra-layer communication overhead. The naive partitioning policy in +Engine overlooks the GPU-CPU intra-layer communication, often offsetting the benefits of assigning high-frequency activation neurons to the GPU. Conversely, our policy in PowerInfer more adeptly balances processing loads and communication costs between the CPU and GPU.
@@
Figure 15: Performance breakdown for each component of PowerInfer on PC-High.
8.3.2 Neuron-aware Operator Performance
@@
Figure 16: Comparing Neuron-aware operator with different sparse operators on PC-Low. The X axis indicates the sparsity level, the Y axis represents the execution time(ms).
@@
This section evaluates the performance of PowerInfer’s sparse operators on both CPU and GPU across various sparsity levels. We benchmark PowerInfer against leading sparse libraries: for CPU benchmarks, we use PyTorch sparse, the state-of-the-art sparse kernels within PyTorch, as our baseline. In GPU, PowerInfer is compared with PIT [51]. Given that the sparsity in LLMs is typically based on neuron granularity, our experiments are specifically designed to evaluate sparse matrices of this nature. We focus on sparse matrix-vector multiplication using a [4096, 4096] $\times$ [4096, 1] configuration, a common setup in local LLM inference [6]. To adjust sparsity, we introduce zero values to matrix rows.
@@
8.3.3 Predictor Overhead
Figure 17: End-to-end prediction overhead of PowerInfer on PC-Low. The X axis represents various models, while the Y-axis displays the percentage breakdown between predictor overhead and LLM inference time.
The execution time of the online predictors for different models is also measured, as depicted in Figure 17. On average, the execution of predictors constitutes less than 10% of the total inference time in PowerInfer. This efficiency is primarily due to the adaptive methods used in constructing sparsity predictors, which minimizes computational load. Moreover, these dense-model predictors are incorporated into PowerInfer’s solver for neuron placement decisions, with a preference for allocating them to the GPU. This strategy effectively leverages the parallel processing capabilities of GPUs, further minimizing the overhead associated with the predictors.
8.3.4 Performance Comparison with A100
Figure 18: Generation speed of NVIDIA RTX 4090 compared with single A100. The X axis represents various models, while the Y-axis represents end-to-end generation speed (tokens/s) under various inference framework. The percentages within the arrows represent the slowdown relative to vLLM on the A100.
In our study, we analyze the extent to which PowerInfer reduces the performance gap between a consumer-grade GPU and its top-tier server-grade counterpart. Therefore, we evaluate the generation speed of PowerInfer, deployed on PC-High, in comparison to the performance of llama.cpp and vLLM [18] executed on a single 80GB NVIDIA A100 GPU. We chose the OPT-30B and Falcon-40B models for comparison, considering their exact memory requirements matching precisely with the capacity of the A100 GPU. Our evaluation used input lengths of 1 and 64 to measure pure generation speed and conversational interactions, respectively.
@@ 8.3.4 Performance Comparison with A100
Table 2: Comparison of LLM accuracy between PowerInfer-optimized models (termed as "model-sparse") and their original counterparts. COPA [34] serves as a QA dataset focused on evaluating causal reasoning skills in language models. PIQA [4] and Winogrande [35] are designed for probing common sense reasoning abilities and the understanding of physical world interactions in LLMs. RTE [15] is used to assess natural language understanding via textual entailment.
PIQA Winogrande RTE COPA
OPT-7B 75.78% 65.19% 55.23% 81.00%
OPT-7B-sparse 75.67% 65.51% 55.96% 81.00%
OPT-13B 76.01% 64.96% 58.12% 85.00%
OPT-13B-sparse 76.28% 65.98% 56.32% 84.00%
OPT-30B 77.58% 68.43% 58.40% 82.00%
OPT-30B-sparse 77.48% 67.56% 59.93% 86.00%
OPT-66B 78.62% 68.82% 60.29% 86.00%
OPT-66B-sparse 79.16% 67.80% 61.01% 87.00%
OPT-175B 79.65% 72.77% 59.93% 88.00%
OPT-175B-sparse 79.26% 72.38% 61.98% 87.00%
Falcon(ReLU)-40B 81.23% 75.45% 66.43% 92.00%
Falcon(ReLU)-40B-sparse 81.01% 75.92% 64.62% 93.00%
LLaMA(ReGLU)-70B 82.01% 75.93% 75.81% 89.00%
LLAMA(ReGLU)-70B-sparse 82.05% 75.53% 77.62% 90.00%
8.4 LLM Accuracy
@@
Since PowerInfer selectively omits neurons predicted to be inactive, we investigated whether this approach affects the inference accuracy of LLMs. Table 2 compares the accuracy of models from the OPT, Falcon (ReLU), and LLaMA (ReGLU) families, both with and without differentiating activated/inactivated neurons, across a variety of downstream tasks. The results show that PowerInfer causes negligible loss in inference accuracy, regardless of the model size or type of task, consistent with previous research findings [21]. Although the predictors in each Transformer layer maintain an accuracy rate above 95%, they may occasionally miss some active neurons. As a result, there are minor fluctuations in LLM accuracy, leading to slight decreases or sometimes even increases in performance on specific downstream tasks.
9 Related Work
LLM Activation Sparsity: Recent advancements like DejaVu [21], PIT [51], and brainstorm [8] are crucial to optimizing LLM inference, akin to PowerInfer. DejaVu [21] proposes enhancing inference speed through activation sparsity prediction, while PowerInfer leverages a power-law distribution in neuron activations, focusing on GPU computation of frequently activated neurons. PIT [51] accelerates GPU tasks by converting sparse to dense matrices. However, these methods, primarily exploiting GPU sparsity, face limitations in resource-constrained local environments.
@@ 9 Related Work
LLM Weight Sparsity: Model pruning [17, 16, 24], reducing parameter count by setting some weights to zero, is exemplified by SparseGPT [11] and Wanda [40], achieving nearly 50% unstructured sparsity. SparTA [52] leverages both sparse tensor and SIMT cores by dividing sparse matrices. Flash-LLM [45] introduces a "Load-as-Sparse and Compute-as-Dense" approach for tensor core SpMM. However, these methods, orthogonal to LLMs’ intrinsic sparse activations, usually incur accuracy losses and wall-clock model acceleration challenges [27]. This is in contrast to the natural sparse activations utilized by PowerInfer, which maintain performance and enhance computational efficiency.
@@
Speculative LLM Inference: speculative inference [13, 6, 43] can also be leveraged to serve models exceeding GPU memory. Speculative decoding [7] uses a smaller, faster model to pre-decode tokens, later validated by the main model in a batch, reducing steps and CPU-GPU communication. SpecInfer [26], as another example, effectively reduces the number of LLM decoding steps and the overall communication between CPU DRAM and GPU HBM. While separate from our focus, integrating speculative inference into PowerInfer could further boost LLM inference speed.
@@
LLM-Specific Serving Optimizations: The prominence of Transformers has led to specialized serving systems [53, 9, 36]. Orca [47] introduces iteration-level scheduling. vLLM [18] implements Paged Attention for token storage in varied GPU memory addresses, overcoming KV cache’s continuous storage limit. While vLLM effectively mitigates the issue of severe GPU memory fragmentation, it does not address the challenge of deploying models on PCs where the entire model cannot fit within the available GPU memory.
10 Conclusion

### v2

@@ 8.3.1. Performance Breakdown
Figure 15 breaks down the contributions of each PowerInfer component to the overall performance speedup. Using a step-by-step integration method, we progressively incorporate PowerInfer features into llama.cpp. First, we add PowerInfer’s predictors and neuron-aware operators into llama.cpp (labeled ”+PO”), enabling computation of only activated neurons on both GPU and CPU. Yet, +PO still adheres to layer-wise computation, where each layer is processed entirely by either GPU or CPU.
Building on +PO, we introduce PowerInfer’s hybrid inference engine (denoted ”+Engine”), which allows neuron-aware operators to process neurons within the same layer simultaneously on both GPU and CPU. +Engine uses a naive neuron partitioning policy that assigns neurons randomly to the GPU. The final step involves integrating our optimized policy (”+Policy”), formulated by the offline solver as described in §$\S$6, into the +Engine setup, showcasing the full capabilities of PowerInfer.
@@
The initial integration of +PO into llama.cpp yields performance boosts of 1.87$\times$ and 3.32$\times$ for Bamboo-7B and Falcon-40B, respectively, primarily by reducing unnecessary inactive neurons. +Engine further escalates these gains to 2.60$\times$ and 7.80$\times$, thanks to precise neuron placement and intra-layer calculations that significantly increase the GPU’s computational share. Finally, incorporating +Policy results in improvements of 3.62$\times$ and 11.69$\times$. The enhancement achieved by our policy lies in its ability to finely balance the intra-layer communication overhead. The naive partitioning policy in +Engine overlooks the hotness of neurons and the GPU-CPU intra-layer communication, often offsetting the benefits of assigning high-frequency activation neurons to the GPU. Conversely, our policy in PowerInfer more adeptly balances processing loads and communication costs between the CPU and GPU.
@@
Figure 15. Performance breakdown for each component of PowerInfer. The Falcon-40B is running on PC-High and Bamboo-7B is running on PC-Low.
8.3.2. Generation Latency Analysis
In this section, we investigate the robustness of PowerInfer’s speedup across different tasks and inter-token generation latency distritbution. We sample from the most representative tasks on the Hugging Face community, including STEM, coding, roleplaying and information extraction. We measure the inter-token generation latency distribution across different tasks. Table 5 presents the evaluation results of Bamboo-7B on PC-Low. PowerInfer exhibits a highly consistent inter-token generation latency across different tasks, with only a 10% difference between the P95 latency and the average. It is due to PowerInfer’s policy that places the general hot neurons on GPU. This fluctuation is caused by the sparsity variation of different tokens from 80% to 86%. When encountering tokens with lower sparsity, the computation introduces 10% more latency compared with the average.
Table 5. Generation latency (ms) distribution on various Tasks.
STEM HumanEval Roleplay Table-Extraction
Bamboo-7B on PC-Low Avg. 38.05 36.71 37.08 37.42
P95 40.75 40.48 40.61 40.74
P99 42.33 42.37 43.87 42.93
8.3.3. Neuron-aware Operator Performance
@@
Figure 16. Comparing Neuron-aware operator with different sparse operators on PC-Low. The X axis indicates the sparsity level, the Y axis represents the execution time (ms).
@@
This section evaluates the performance of PowerInfer’s sparse operators on both CPU and GPU across various sparsity levels. We benchmark PowerInfer against leading sparse libraries: for CPU benchmarks, we use PyTorch sparse, the state-of-the-art sparse kernels within PyTorch, as our baseline. In GPU, PowerInfer is compared with PIT (Zheng et al., 2023b). Given that the sparsity in LLMs is typically based on neuron granularity, our experiments are specifically designed to evaluate sparse matrices of this nature. We focus on sparse matrix-vector multiplication using a [4096, 4096] $\times$ [4096, 1] configuration, a common setup in local LLM inference (Cai et al., 2023). To adjust sparsity, we introduce zero values to matrix rows.
@@
8.3.4. Predictor Overhead
Figure 17. End-to-end prediction overhead of PowerInfer on PC-Low. The Y-axis displays the percentage breakdown between predictor overhead and LLM inference time.
Table 6. Predictor parameter sizes and ratios for various LLMs. The predictor ratio represents the percentage of predictor parameters relative to the original model parameters.
Model OPT-13B OPT-66B Falcon(ReLU)-40B LLaMA(ReGLU)-70B
Predictor-params 0.88B 3.23B 3.63B 5.66B
Predictor-ratio(%) 6.71% 4.88% 8.68% 8.08%
The execution time of the online predictors for different models is also measured, as depicted in Figure 17. On average, the execution of predictors constitutes less than 10% of the total inference time in PowerInfer. The execution time of predictors correlates directly with the predictor sizes, which, as shown in Table 6, comprise only 7.09% of the model weights. The efficiency of these predictors stems from adaptive construction methods that minimize both size and computational load. Furthermore, PowerInfer integrates these predictors into its solver for neuron placement decisions, preferentially allocating them to GPUs. This strategy exploits the parallel processing capabilities of GPUs, further reducing the runtime overhead associated with prediction.
8.3.5. Performance Comparison with A100
Figure 18. Generation speed of NVIDIA RTX 4090 compared with single A100. The X axis represents various models, while the Y axis represents generation speed (tokens/s) under various inference framework. The percentages within the arrows represent the slowdown relative to vLLM on the A100.
In our study, we analyze the extent to which PowerInfer reduces the performance gap between a consumer-grade GPU and its top-tier server-grade counterpart. Therefore, we evaluate the generation speed of PowerInfer, deployed on PC-High, in comparison to the performance of llama.cpp and vLLM (Kwon et al., 2023) executed on a single 80GB NVIDIA A100 GPU. We chose the OPT-30B and Falcon-40B models for comparison, considering their exact memory requirements matching precisely with the capacity of the A100 GPU. Our evaluation used input lengths of 1 and 64 to measure pure generation speed and conversational interactions, respectively.
@@ 8.3.5. Performance Comparison with A100
Table 7. Comparison of LLM accuracy between PowerInfer-optimized models (termed as ”model-PowerInfer”) and their original counterparts. Arc-Challenge (Clark et al., 2018) is a dataset for evaluating AI systems’ comprehension and reasoning in natural language. MMLU (Hendrycks et al., 2021) benchmarks model performance across various domains. PIQA (Bisk et al., 2020) and Winogrande (Sakaguchi et al., 2021) assess common sense reasoning and understanding of physical interactions in LLMs.
PIQA Winogrande Arc-Challenge MMLU GSM8K Average
OPT-7B 75.78% 65.19% 30.63% 24.95% 1.90% 39.69%
OPT-7B-PowerInfer 75.67% 65.51% 30.63% 24.73% 1.82% 39.67%
OPT-13B 76.01% 64.96% 32.94% 25.02% 2.12% 40.21%
OPT-13B-PowerInfer 76.28% 65.98% 33.19% 24.76% 2.20% 40.48%
LLaMA(ReGLU)-13B 76.44% 70.09% 36.52% 50.21% 25.40% 51.73%
LLaMA(ReGLU)-13B-PowerInfer 74.06% 69.93% 36.60% 49.47% 23.90% 50.79%
Falcon-40B 81.23% 75.45% 50.68% 51.78% 21.99% 56.23%
Falcon-40B-PowerInfer 81.01% 75.92% 50.68% 51.68% 20.45% 55.95%
LLaMA(ReGLU))-70B 82.01% 75.93% 52.39% 62.30% 62.30% 66.99%
LLAMA(ReGLU)-70B-PowerInfer 82.05% 75.53% 51.45% 61.90% 61.90% 66.57%
8.4. SiLU-based LLM Performance
Table 8. Generation speed (tokens/s) comparison for SwiGLU and ReLU-based LLMs.
Setting Model PowerInfer llama.cpp Speedup
PC-High Yi(SwiGLU)-34B 1.7 1.0 1.7$\times$
PC-Low LLaMA(SwiGLU)-2-13B 3.1 2.1 1.5$\times$
PC-High OPT-30B 12.0 1.1 10.9$\times$
While PowerInfer demonstrates remarkable speedups on ReLU-based models due to their high activation sparsity, it also shows effectiveness in accelerating SiLU-based models. Our evaluation of SiLU-based LLMs, as presented in Table 8, reveals that PowerInfer achieves a speedup of 1.47$\times$ to 1.7$\times$. This performance gain, albeit less pronounced than for ReLU-based counterparts (like OPT-30B in Table 8), underscores the efficacy of PowerInfer’s sparse computation mechanisms across different activation functions. The comparatively lower speedup can be attributed to the reduced sparsity in SiLU-based models, where a larger proportion of neurons remain active during inference. Consequently, CPU computation emerges as a potential bottleneck in these scenarios. The speedup observed in SiLU models indicates that the effectiveness of PowerInfer’s acceleration is correlated with the extent of the model’s sparsity.
8.5. LLM Accuracy
@@
Since PowerInfer selectively omits neurons predicted to be inactive, we investigated whether this approach affects the inference accuracy of LLMs, focusing on models using ReLU or related activation functions. Table 7 compares the accuracy of models from the OPT, Falcon (ReLU), and LLaMA2 (ReGLU) families, both with and without differentiating activated/inactivated neurons, across a variety of representative downstream tasks. The results show that PowerInfer causes negligible loss in inference accuracy, regardless of the model size or type of task, consistent with previous research findings (Liu et al., 2023b). Although the predictors in each Transformer layer maintain an accuracy rate above 95%, they may occasionally miss some active neurons. As a result, there are minor fluctuations in LLM accuracy, leading to slight fluctuations in performance on specific downstream tasks.
We further analyze the nature and impact of mispredicted neurons on inference accuracy. Our investigation reveals that neurons mispredicted by the predictor typically have minimal influence on the layer’s output. This characteristic makes these neurons more susceptible to misprediction, as their activation or non-activation has little effect on the overall result. To quantify this impact, we conducted an analysis on the OPT-7B model. By comparing the cosine similarity between outputs from original dense computations and our predictor-based computations, we found that mispredictions only affect results by approximately 0.4%. This minimal difference further supports our observation that PowerInfer maintains the accuracy of the original model, despite the occasional misprediction of neurons.
Table 9. Performance comparison between SLM (Qwen-1.5-4B) and PowerInfer (Bamboo-7B).
Model TBT(ms) Average(%) MMLU(%) GSM8K(%) ARC-C(%)
Qwen1.5-4B 10.83 52.23 55.26 53.9 47.53
Bamboo-7B-PowerInfer 11.85 65.65 62.26 70.54 64.16
Bamboo-7B-dense 18.54 65.49 62.46 70.28 63.74
Comparison with Smaller Language Models. To better understand PowerInfer’s performance gains, we compare it with a common acceleration approach: using smaller language models (SLMs). However, the SLM approach often comes at the cost of reduced model accuracy. To fully understand the benefits of PowerInfer, we compare its performance using a larger model to that of a state-of-the-art SLM, evaluating both speed and accuracy. Table 9 shows that the Bamboo-7B model outperforms the Qwen-1.5-4B (a state-of-the-art 4B model) across various tasks, and PowerInfer (Bamboo-7B-PowerInfer) maintains the original dense model’s accuracy while achieving 4B-level decoding speed.
9. Related Work
@@
LLM Weight Sparsity: Model pruning (Han et al., 2015b, a; Ma et al., 2023b) reduces parameters by setting weights to zero, as seen in SparseGPT (Frantar and Alistarh, 2023) and Wanda (Sun et al., 2023), achieving 50% sparsity. SparTA (Zheng et al., 2022) leverages both sparse tensor and SIMT cores by dividing sparse matrices. Flash-LLM (Xia et al., 2023) introduces a ”Load-as-Sparse and Compute-as-Dense” approach for tensor core SpMM. However, these methods, orthogonal to LLMs’ intrinsic sparse activations, usually incur accuracy losses and wall-clock model acceleration challenges (Mirzadeh et al., 2023). In contrast, PowerInfer uses natural sparse activations to maintain performance and efficiency.
LLM Attention Sparsity: PowerInfer leverages the sparse activation characteristics of MLP layers. It’s worth noting that attention blocks also exhibit sparsity, which primarily stems from two sources: attention heads, where only some need computation for the current token (Liu et al., 2023b), and sparsity within heads, where unimportant KV cache is pruned or offloaded to alleviate memory-bound bottleneck of LLMs serving, as in H2o (Zhang et al., 2024a) and InfiniGen (Lee et al., 2024a). The sparse activations used by PowerInfer are orthogonal to attention sparsity and can further reduce inference latency.
@@
Speculative LLM Inference: Speculative inference (Fu et al., 2023; Cai et al., 2023; Wang et al., 2023; Chen et al., 2023a) can also be leveraged to serve models exceeding GPU memory, which uses a smaller model to pre-decode tokens, later validated by the main model in a batch. SpecInfer (Miao et al., 2023) effectively reduces the number of LLM decoding steps. While separate from our focus, integrating speculative inference into PowerInfer could further boost LLM inference speed.
@@
LLM-Specific Serving Optimizations: There are many ML serving systems (Gujarati et al., 2020) for general ML models. The prominence of Transformers has led to specialized serving systems (Zhou et al., 2022; Fang et al., 2021; Sheng et al., 2023a; Chen et al., 2023b; Khare et al., 2023). Orca (Yu et al., 2022) introduces iteration-level scheduling. vLLM (Kwon et al., 2023) implements PagedAttention for token storage in varied GPU memory addresses, overcoming KV cache’s continuous storage limit. These methods do not address the challenge of deploying models on PCs where the entire model cannot fit within the GPU memory.
10. Conclusion
@@
11. Acknowledgments
We sincerely thank our shepherd Shivaram Venkataraman and anonymous reviewers for their insightful suggestions. We are deeply grateful to Rong Chen and Yubin Xia for their comprehensive and constructive feedback, which greatly enhanced the quality of this paper. This work was partially supported by National Key R&D Program of China (2023YFB4503702) and NSFC (No. 62372287 and 61925206). Zeyu Mi (yzmizeyu@sjtu.edu.cn) is the corresponding author.